# Turán number of complete multipartite graphs in multipartite graphs

### v1

@@ 4. Proof of the upper bounds
Now we prove the upper bounds on $\operatorname{ex}_{k}(n,K_{r}(t))$.
The proof of Theorem 6 is done by a series of claims and estimates, and actually contains the proofs of Theorems 5 and 7. More precisely, the first half of the proof of Theorem 6 only requires ($Z$) and (consequently) ($E1$), but not ($E2$) or ($E3$). The (weaker) estimates obtained in this part of the proof suffice for Theorems 5 and 7. The second part of the proof contains refined estimates obtained from ($E2$) and ($E3$) and completes the proof of Theorem 6.
Outline of the proofs. Now we give an outline of our proofs. Let $G\in\mathcal{G}_{k}(n)$ be $K_{r+1}(t)$-free and has the maximum number of edges. Since $e(G)>t_{r}(k)n^{2}$, we can assume that $G$ is $\gamma$-close to some $T\in{\mathcal{T}}_{r,k}(n)$. Using that $e(G)$ is maximum, we can easily derive a minimum degree condition by some symmetrization arguments.
Next we define atypical vertices. Roughly speaking, there are two types of atypical vertices: the first type of vertices, denoted by $Z^{\prime\prime}\cup W^{\prime\prime}$, are the “wrong” ones that do not exist in ${\mathcal{T}}_{r,k}(n)$; the second type of vertices, denoted by $(W^{\prime}_{i}\setminus W_{i})\cup\bigcup_{j\neq i}Z_{j}^{i}$ for $i\in[r]$, are the vertices that are not in $U_{i}$ but behave like the vertices of $U_{i}$, in other words, they are in the wrong place. In the first half of the proof (i.e., the proofs of Theorem 5 and Theorem 7), we completely ignore the first type of atypical vertices because there are only a constant number of them (see $(P1)$ and $(P3)$) and they contribute only $O(n)$ to $e(G)$. For the second type of atypical vertices, there are only $o(n)$ of them (see $(P1)$ and $(P3)$) and we move them to appropriate rows and redefine our partition as $\tilde{U}_{1},\dots,\tilde{U}_{r}$ (see (4.2)). A key observation is that $Z_{j}^{i}\neq\emptyset$ (namely, there is a vertex in $Z_{j}$ but behaves as a vertex in $Z_{i}$) is possible only if $|W_{j}|\geq(1-o(1))n$.
Now we estimate $e(G)$. We split $E(G)$ into $E_{G}(\tilde{U}_{1}),\dots,E_{G}(\tilde{U}_{r})$, and $E(G^{\prime})$, where $G^{\prime}:=G\cap K(\tilde{U}_{1},\dots,\tilde{U}_{r})$. We have a relatively good estimate of $e(G^{\prime})$ (see Claim 4.4) taking into account that the partition is no longer balanced. In contrast, due to the second type of atypical vertices, we can only show that each $G[\tilde{U}_{i}]$ is “almost” $K_{t,t}$-free (see Claim 4.5). Similarly, we show that all but at most one rows are “almost” $K_{1,t}$-free (see Claim 4.7). Assuming that $e_{G}(\tilde{U}_{1})$ is the largest among all $e_{G}(\tilde{U}_{i})$, $i\in[r]$, we can collect the estimates and give an upper bound of $e(G)$. Now we show that $\tilde{U}_{1}$ has no atypical vertices (Claim 4.8), and thus $e_{G}(\tilde{U}_{1})\leq z_{t}^{(a+1)}(n)$. We further refine our estimate on $\tilde{U}_{i}$, $i>1$, and show that each second type atypical vertex contributes at most a constant number of edges to $E(G)\setminus E(G^{\prime})$ (Claims 4.9 and 4.10). In summary, $\tilde{U}_{1}$ is indeed $K_{t,t}$-free, $e_{G}(\tilde{U}_{i})=O(n)$ for $i>1$, and $|Z^{\prime\prime}\cup W^{\prime\prime}|=O(1)$, from which we conclude the proofs of Theorem 5 and Theorem 7.
To prove Theorem 6, we refine earlier estimates as follows. We first show that $|W_{1}^{\prime}|=(1-o(1))n$ and $Z^{\prime\prime}\cup W^{\prime\prime}=\emptyset$, where we use ($E2$). The rest of the proofs are further refinements of our estimates for various cases (see, e.g. the definition of $h_{2}$). In particular, at the very last step, we use ($E3$) to show that all of $Z_{1}^{i}$, $i\in[2,r]$ should be empty and $|W_{1}^{\prime}|$ should be equal to $n$.
We start with the following simple proposition.
@@
Given integers $1\leq s\leq t$ and sufficiently large $m,n$, Kővári, Sós, Turán [MR0065617] showed that $z(m,n,s,t)\leq Cmn^{1-1/s}$ for some $C=C(t)>0$, that is, a bipartite graph $G$ with parts of size $m$ and $n$ has at most $Cmn^{1-1/s}$ edges if $G$ has no copy of $K_{s,t}$ where the part of size $s$ is in the part of $G$ of size $m$. This easily implies that $\operatorname{ex}(n,K_{s,t})\leq Cn^{2-1/t}$ for sufficiently large $n$.
Proofs of Theorems 5, 6, and 7.
Suppose ($Z$) holds, that is, $z_{t}^{(2)}(n)\geq cn^{2-1/t}$ for some $c>0$. Take $C=C(t)$ as in the Kővári–Sós–Turán result in the previous paragraph. We choose constants
@@
Suppose $G$ is $K_{r+1}(t)$-free and has the maximum number of edges, that is, $e(G)=\operatorname{ex}_{k}(n,K_{r+1}(t))$. Suppose further that $e(G)>g_{2}(n,r,k,t)>t_{r}(k)n^{2}$. By Theorem 3, $G$ is $\gamma$-close to some $T\in{\mathcal{T}}_{r,k}(n)$, where the color classes of $T$ are denoted by $U_{1},\dots,U_{r}$ with $U_{i}=W_{i}\cup Z_{i}$ such that $Z_{i}=V_{(i-1)a+1}\cup\cdots\cup V_{ia}$, $W_{i}=\emptyset$ if $b=0$ and $W_{i}$ is a subset of $V_{j}$ for some $j\geq ar+1$ otherwise, and $T=K(U_{1},\dots,U_{r})\cap K(V_{1},\dots,V_{k})$. For $i\in[r]$, if $W_{i}\neq\emptyset$, then let $q_{i}$ be the index such that $W_{i}\subseteq V_{q_{i}}$ and we know $ar<q_{i}\leq k$. For simplicity, we write $z_{t}(n)=z_{t}^{(\lceil k/r\rceil)}(n)$.
@@
Let $Z_{i}^{\prime}:=Z_{i}\setminus Z_{i}^{\prime\prime}$ for all $i$. We write $Z_{i}^{\prime}$ as $\bigcup_{j\in[r]}Z_{i}^{j}$, where $Z_{i}^{j}$, $j\neq i$, consists of the vertices $v$ such that $d(v,Z_{j})<\varepsilon n$, and $Z_{i}^{i}$ consists of the vertices $v$ such that $d(v,U_{i})<\varepsilon n$. Below are some useful properties of these sets.
@@ 3. $(Q3)$
Estimate $e(G)$
Let $G^{\prime}=G\cap K(\tilde{U}_{1},\dots,\tilde{U}_{r})$. We have $e(G)=e(G^{\prime})+\sum_{i=1}^{r}e_{G}(\tilde{U}_{i})+e(Z^{\prime\prime}\cup W^{\prime\prime},G)$.
Since $G^{\prime}$ is $r$-partite, it is $K_{r+1}$-free. As $G^{\prime}$ is a subgraph of $G\in\mathcal{G}_{k}(n)$, we have $e(G^{\prime})\leq t_{r}(k)n^{2}$ (but this is not good enough when $b>0$). Below we give an upper bound for $e(G^{\prime})$, which will be used throughout the proof. Recall that $T=K(V_{1},\dots,V_{k})\cap K(U_{1},\dots,U_{r})$ has precisely $t_{r}(k)n^{2}$ edges,
@@
Our main task is bounding the number of edges in each $\tilde{U}_{i}$. For $i\in[r]$, we have $e(G[\tilde{U}_{i}])=e(Z_{i}^{i},G[\tilde{U}_{i}])+e_{G}(\tilde{U}_{i}\setminus Z_{i}^{i})$. To bound $e_{G}(\tilde{U}_{i}\setminus Z_{i}^{i})=e_{G}((\tilde{Z}_{i}\setminus Z_{i})\cup W_{i}^{\prime})$, we note that $e_{G}(\tilde{Z}_{i}\setminus Z_{i},W_{i}^{\prime})\leq|\tilde{Z}_{i}\setminus Z_{i}||W_{i}^{\prime}|$ and $e_{G}(W_{i}^{\prime})\leq e_{T}(W_{i}^{\prime})$. However, we may not have $e_{G}(\tilde{Z}_{i}\setminus Z_{i})\leq e_{T}(\tilde{Z}_{i}\setminus Z_{i})$ because each $Z_{j}^{i}$ is an independent set in $T$, but may not be independent in $G$ when $a\geq 2$. Thus, $e_{G}(\tilde{Z}_{i}\setminus Z_{i})\leq e_{T}(\tilde{Z}_{i}\setminus Z_{i})+\sum_{j\neq i}e_{G}(Z_{j}^{i})$. Putting these together, for each $i\in[r]$, we have
@@
Let $f_{i}:=e(Z_{i}^{i},G[\tilde{U}_{i}])$. Applying Claim 4.4 and (4.5), we derive that
$e(G)$ $=e(G^{\prime})+e(Z^{\prime\prime}\cup W^{\prime\prime},G)+\sum_{i\in[r]}f_{i}+e_{G}((\tilde{U}_{i}\setminus Z^{i}_{i})$ (4.6)
@@
The following claim shows that a large portion of $G[\tilde{U}_{i}]$ is $K_{t,t}$-free (though we cannot prove that the entire $G[\tilde{U}_{i}]$ is $K_{t,t}$-free).
@@ 3. $(K3)$
We now derive an upper bound for $e(G)$ from Claims 4.4 and 4.5. For $i\in[r]$, we have $\beta_{i}\leq\sum_{j\in L\setminus\{i\}}|Z_{j}^{i}|(|\tilde{Z}_{j}\setminus Z_{j}|+|W_{j}^{\prime}\setminus V_{q_{j}}|+|Z_{i}\setminus\tilde{Z}_{i}|)$ as $|W_{j}^{\prime}|-n\leq|W_{j}^{\prime}\setminus V_{q_{j}}|$. Fix $j\in L\setminus\{i\}$. Note that $|W_{j}|\geq(1-2\varepsilon)n$. We have $|\tilde{Z}_{j}\setminus Z_{j}|\leq t-1$ by $(Q3)$, and $|W_{j}^{\prime}\setminus V_{q_{j}}|\leq t-1$ by $(K2)$. If $|W_{i}|>n/2$, then $|Z_{j}^{i}|\leq t-1$ by $(Q3)$. Furthermore, since $|Z_{i}\setminus\tilde{Z}_{i}|\leq(r-1)\sqrt{\gamma}n+C_{0}$ by $(P3)$, it follows that
@@
$e(G)$ $\leq t_{r}(k)n^{2}+(r+1)C_{0}kn+\sum_{i\in[r]}f_{i}+2(r-1)rC_{0}\sqrt{\gamma}n+\sum_{i,j:i\neq j}e_{G}(Z_{j}^{i})$ (4.9)
$\leq t_{r}(k)n^{2}+\sum_{i\in[r]}f_{i}+r^{2}C\sqrt{\gamma}n^{2-1/t}.$
Using the assumption $e(G)\geq g_{2}(n,r,k,t)\geq t_{r}(k)n^{2}+z_{t}(n)$, we infer that
@@
In the rest of the proof we will derive a contradiction to (4.10). We first study the existence of $K_{1,t}$ in each color class. To do so, we consider a copy of $K_{3}(t)$ in $G[\tilde{U}_{i}\cup\tilde{U}_{j}]$ for some $i\neq j$.
@@
First assume that there exists $j\in[r]$ such that $G[\tilde{U}_{j}]$ contains a copy of $K_{1,t}$, with vertex set denoted by $\{v,u_{1},\dots,u_{t}\}$, $v\in\tilde{U}_{j}$ and $u_{1},\dots,u_{t}\in Z_{j}^{j}$. Fix $i\in[r]\setminus\{j\}$ and let $N^{\prime}$ be the set of common neighbors of $u_{1},\dots,u_{t}$ in $\tilde{U}_{i}\cap U_{i}$. Suppose $v\in V_{p}$ and let $N$ be the set of common neighbors of these $t+1$ vertices in $\tilde{U}_{i}\cap U_{i}$. In particular, $N\subseteq N^{\prime}$ and $N$ is almost equal to the union of $a$ or $a+1$ clusters in $\tilde{U}_{i}$. Suppose there is a copy of $K_{t-1,t}$ with parts $S_{1}$ of size $t-1$ and $S_{2}$ of size $t$ such that $S_{1}\subseteq N^{\prime}$ and $S_{2}\subseteq N$. Then by Claim 4.6, there exists $l\in[r]\setminus\{i,j\}$ such that $B\cap Z_{l}\neq\emptyset$ and $B\cap V_{q_{l}}\neq\emptyset$, where $B$ denotes the vertex set of the copy of $K_{3}{(t)}$. This is impossible since $v$ is the only possible vertex in $B\cap(Z_{l}\cup V_{q_{l}})$ and can not satisfy both. Therefore, $G[N,N^{\prime}]$ is $K_{t-1,t}$-free, implying that $e_{G}(N,N^{\prime})=O(n^{2-1/(t-1)})$.
By $(P1)$$(P3)$ and (Deg), we have $|(\tilde{U}_{i}\setminus V_{p})\setminus N|\leq 3(t+1)\varepsilon n$ and $|\tilde{U}_{i}\setminus N^{\prime}|\leq 3t\varepsilon n$. Let $E^{i}$ be the set of the edges incident to $(\tilde{U}_{i}\setminus V_{p})\setminus N$ or $\tilde{U}_{i}\setminus N^{\prime}$ and counted in $f_{i}$. We split it to $E^{i}\cap E_{G}(Z_{i}^{i})$ and $E^{i}\cap E_{G}(\tilde{U}_{i}\setminus Z_{i}^{i},Z_{i}^{i})$. Note that by $(K1)$, each of the terms can be split further into at most $k$ $K_{t,t}$-free bipartite graphs, each with one part of size at most $3(t+1)\varepsilon n$ and the other part of size at most $(1+(r-2)\sqrt{\gamma})an$. Therefore, we obtain that
@@
for $\text{for }i\geq 2,\quad d(v,Z_{i}^{i})\leq t-1\ \text{for all}\ v\in\tilde{U}_{i},\quad\text{and thus},\quad f_{i}\leq(t-1)|\tilde{U}_{i}|.$ (4.12)
If $b=0$, then $\tilde{U}_{i}=Z_{i}^{i}=Z_{i}\setminus Z^{\prime\prime}$ for all $i$ by (4.3). In this case $\tilde{U}_{1}$ is $K_{t,t}$-free by $(K1)$ and $\tilde{U}_{i}$ is $K_{1,t}$-free for all $i\geq 2$ by (4.12). Since $G$ is $\gamma$-close to $K_{r}(an)$, $G[U_{i}\setminus Z^{\prime\prime},U_{j}\setminus Z^{\prime\prime}]$ is almost complete for all $i\neq j$. This completes the proof of Theorem 7 with $Z:=Z^{\prime\prime}$ (note that ($Z$) holds because $t=2,3$).
Furthermore, when $b=0$, together with $e(G^{\prime})\leq t_{r}(k)n^{2}$, we conclude that
$e(G)\leq t_{r}(k)n^{2}+z_{t}(n)+(r-1)\left\lfloor\frac{t-1}{2}an\right\rfloor+|Z^{\prime\prime}|kn.$ (4.13)
By $(P3)$, $|Z^{\prime\prime}|\leq rC_{0}$, and this proves Theorem 5 for the case $b=0$.
We thus assume $b>0$ for the rest of the proof of Theorem 5. By (4.10) and (4.12), we get
$f_{1}\geq z_{t}^{(a+1)}(n)-\varepsilon n^{2-1/t}.$ (4.14)
@@
Now we can give a much cleaner structure, shown in a series of claims below.

### v2

@@ 3.3. Two quick proofs
We need the following simple proposition.
@@
4. Main Proofs
Given integers $1\leq s\leq t$ and sufficiently large $m,n$, Kővári, Sós, Turán [MR0065617] showed that $z(m,n,s,t)\leq Cmn^{1-1/s}$ for some $C=C(t)>0$, that is, a bipartite graph $G$ with parts of size $m$ and $n$ has at most $Cmn^{1-1/s}$ edges if $G$ has no copy of $K_{s,t}$ where the part of size $s$ is in the part of $G$ of size $m$. We start the proof with the general setting, that is, for $k=ar+b$ with $0\leq b<r$. After we conclude the proof of Theorem 4, we focus on the case $k\leq 2r$.
Proofs of Theorems 3 and 4.
Suppose $t=2,3$ and thus ($Z$) holds, that is, $z_{t}^{(2)}(n)\geq cn^{2-1/t}$ for some $c>0$. Take $C=C(t)$ as in the Kővári–Sós–Turán result in the previous paragraph. We choose constants
@@ 4. Main Proofs
Suppose $G$ is $K_{r+1}(t)$-free and has the maximum number of edges, that is, $e(G)=\operatorname{ex}_{k}(n,K_{r+1}(t))$. Suppose further that $e(G)>g(n,r,k,t)>t_{r}(k)n^{2}$. By Theorem 5, $G$ is $\gamma$-close to some $T\in{\mathcal{T}}_{r,k}(n)$ such that $T=K(U_{1},\dots,U_{r})\cap K(V_{1},\dots,V_{k})$, where $U_{1},\dots,U_{r}$ is a partition of $V(T)$ satisfying the following properties:
  • $U_{i}=W_{i}\cup Z_{i}$ such that $Z_{i}=V_{(i-1)a+1}\cup\cdots\cup V_{ia}$, and
  • $W_{i}=\emptyset$ if $b=0$ and $W_{i}$ is a subset of $V_{q_{i}}$ for some $q_{i}$ with $ar<q_{i}\leq k$ otherwise.
For simplicity, we write $z_{t}(n)=z_{t}^{(\lceil k/r\rceil)}(n)$.
@@
Furthermore, let $Z_{i}^{\prime}:=Z_{i}\setminus Z_{i}^{\prime\prime}$ and write $Z_{i}^{\prime}$ as $\bigcup_{j\in[r]}Z_{i}^{j}$, where $Z_{i}^{j}$, $j\neq i$, consists of the vertices $v\in Z_{i}$ such that $d(v,Z_{j})<\varepsilon n$, and $Z_{i}^{i}$ consists of the vertices $v$ such that $d(v,U_{i})<\varepsilon n$. The following are some useful properties of these sets.
@@ 3. $(Q3)$
First Estimate on $e(G)$
Let $G^{\prime}=G\cap K(\tilde{U}_{1},\dots,\tilde{U}_{r})$. We have $e(G)=e(G^{\prime})+\sum_{i=1}^{r}e_{G}(\tilde{U}_{i})+e(Z^{\prime\prime}\cup W^{\prime\prime},G)$. Since $G^{\prime}$ is $r$-partite, it is $K_{r+1}$-free. As $G^{\prime}$ is a subgraph of $G\in\mathcal{G}_{k}(n)$, we have $e(G^{\prime})\leq t_{r}(k)n^{2}$ (but this is not good enough when $b>0$). Below we give an upper bound for $e(G^{\prime})$, which will be used throughout the proof. Recall that $T=K(V_{1},\dots,V_{k})\cap K(U_{1},\dots,U_{r})$ has precisely $t_{r}(k)n^{2}$ edges.
@@
What remains is to estimate the number of edges in each $\tilde{U}_{i}$. For $i\in[r]$, we have $e(G[\tilde{U}_{i}])=e(Z_{i}^{i},G[\tilde{U}_{i}])+e_{G}(\tilde{U}_{i}\setminus Z_{i}^{i})$. To bound $e_{G}(\tilde{U}_{i}\setminus Z_{i}^{i})=e_{G}((\tilde{Z}_{i}\setminus Z_{i})\cup W_{i}^{\prime})$, we note that $e_{G}(\tilde{Z}_{i}\setminus Z_{i},W_{i}^{\prime})\leq|\tilde{Z}_{i}\setminus Z_{i}||W_{i}^{\prime}|$ and $e_{G}(W_{i}^{\prime})\leq e_{T}(W_{i}^{\prime})$. However, we may not have $e_{G}(\tilde{Z}_{i}\setminus Z_{i})\leq e_{T}(\tilde{Z}_{i}\setminus Z_{i})$ because each $Z_{j}^{i}$ is an independent set in $T$, but may not be independent in $G$ when $a\geq 2$. Thus, $e_{G}(\tilde{Z}_{i}\setminus Z_{i})\leq e_{T}(\tilde{Z}_{i}\setminus Z_{i})+\sum_{j\neq i}e_{G}(Z_{j}^{i})$. Putting these together, for each $i\in[r]$, we have
@@
Let $f_{i}:=e(Z_{i}^{i},G[\tilde{U}_{i}])$. By Claim 4.3, (4.5) and $e(G)=e(G^{\prime})+e(Z^{\prime\prime}\cup W^{\prime\prime},G)+\sum_{i\in[r]}e_{G}(\tilde{U}_{i})$, we derive that
$e(G)$ $\leq t_{r}(k)n^{2}+e(Z^{\prime\prime}\cup W^{\prime\prime},G)+\sum_{i\in[r]}\left(f_{i}+\beta_{i}-\alpha_{i}+e_{G}(\tilde{U}_{i}\setminus Z_{i}^{i})\right)$ (4.6)
@@
We now focus on the structure of each $\tilde{U}_{i}$. We first show that $G[\tilde{U}_{i}]$ is “almost” $K_{t,t}$-free.
@@ 3. $(K3)$
We now derive a lower bound for $\sum f_{i}$ from Claims 4.3 and 4.4. For $i\in[r]$, we have $\beta_{i}\leq\sum_{j\in L\setminus\{i\}}|Z_{j}^{i}|(|\tilde{Z}_{j}\setminus Z_{j}|+|W_{j}^{\prime}\setminus V_{q_{j}}|+|Z_{i}\setminus\tilde{Z}_{i}|)$ as $|W_{j}^{\prime}|-n\leq|W_{j}^{\prime}\setminus V_{q_{j}}|$. Fix $j\in L\setminus\{i\}$. Note that $|W_{j}|\geq(1-2\varepsilon)n$. We have $|\tilde{Z}_{j}\setminus Z_{j}|\leq t-1$ by $(Q3)$, and $|W_{j}^{\prime}\setminus V_{q_{j}}|\leq t-1$ by $(K2)$. If $|W_{i}|>n/2$, then $|Z_{j}^{i}|\leq t-1$ by $(Q3)$. Furthermore, since $|Z_{i}\setminus\tilde{Z}_{i}|\leq(r-1)\sqrt{\gamma}n+C_{0}$ by $(P3)$, it follows that
@@
$e(G)$ $\leq t_{r}(k)n^{2}+(r+1)C_{0}kn+\sum_{i\in[r]}f_{i}+2(r-1)rC_{0}\sqrt{\gamma}n+\sum_{i,j:i\neq j}e_{G}(Z_{j}^{i})$
$\leq t_{r}(k)n^{2}+\sum_{i\in[r]}f_{i}+r^{2}C\sqrt{\gamma}n^{2-1/t},$
as $\gamma\ll 1$. Using the assumption $e(G)\geq g(n,r,k,t)\geq t_{r}(k)n^{2}+z_{t}(n)$, we infer that
@@
We next study the existence of $K_{1,t}$ in each color class. To do so, we consider a copy of $K_{3}(t)$ in $G[\tilde{U}_{i}\cup\tilde{U}_{j}]$ for some $i\neq j$.
@@
First assume that there exists $j\in[r]$ such that $G[\tilde{U}_{j}]$ contains a copy of $K_{1,t}$, with vertex set denoted by $\{v,u_{1},\dots,u_{t}\}$, $v\in\tilde{U}_{j}$ and $u_{1},\dots,u_{t}\in Z_{j}^{j}$. Fix $i\in[r]\setminus\{j\}$ and let $N^{\prime}$ be the set of common neighbors of $u_{1},\dots,u_{t}$ in $\tilde{U}_{i}\cap U_{i}$. Suppose $v\in V_{p}$ and let $N$ be the set of common neighbors of these $t+1$ vertices in $\tilde{U}_{i}\cap U_{i}$. In particular, $N\subseteq N^{\prime}$ and $N$ is almost equal to the union of $a$ or $a+1$ clusters in $\tilde{U}_{i}$. Suppose there is a copy of $K_{t-1,t}$ with parts $S_{1}$ of size $t-1$ and $S_{2}$ of size $t$ such that $S_{1}\subseteq N^{\prime}$ and $S_{2}\subseteq N$. Then by Claim 4.5, there exists $l\in[r]\setminus\{i,j\}$ such that $B\cap Z_{l}\neq\emptyset$ and $B\cap V_{q_{l}}\neq\emptyset$, where $B$ denotes the vertex set of the copy of $K_{3}{(t)}$. This is impossible since $v$ is the only possible vertex in $B\cap(Z_{l}\cup V_{q_{l}})$ and can not satisfy both. Therefore, letting $N^{*}=N\cup(N^{\prime}\cap V_{p})$, we infer that $e_{G}(N^{*})=e_{G}(N)+e_{G}(N,N^{\prime}\setminus N)=O(n^{2-1/(t-1)})$.
By $(P1)$$(P3)$ and (Deg), we have $|\tilde{U}_{i}\setminus N^{*}|\leq 3(t+1)\varepsilon n$. Let $E^{i}$ be the set of the edges incident to $\tilde{U}_{i}\setminus N^{*}$ and counted in $f_{i}$. We split it to $E^{i}\cap E_{G}(Z_{i}^{i})$ and $E^{i}\cap E_{G}(\tilde{U}_{i}\setminus Z_{i}^{i},Z_{i}^{i})$. Note that by $(K1)$, each of the terms can be split further into at most $k$ $K_{t,t}$-free bipartite graphs, each with one part of size at most $3(t+1)\varepsilon n$ and the other part of size at most $(1+(r-2)\sqrt{\gamma})an$. Therefore, we obtain that
@@
for $\text{for }i\geq 2,\quad d(v,Z_{i}^{i})\leq t-1\ \text{for all}\ v\in\tilde{U}_{i},\quad\text{and thus},\quad f_{i}\leq\begin{cases}(t-1)|\tilde{U}_{i}|\hfill\text{ if }a\geq 2,\\ (t-1)|\tilde{U}_{i}\setminus Z_{i}^{i}|\hfill\text{ if }a=1.\end{cases}$ (4.11)
If $b=0$, then $\tilde{U}_{i}=Z_{i}^{i}=Z_{i}\setminus Z^{\prime\prime}$ for all $i$ by (4.3). In this case $\tilde{U}_{1}$ is $K_{t,t}$-free by $(K1)$ and $\tilde{U}_{i}$ is $K_{1,t}$-free for all $i\geq 2$ by (4.11). Since $G$ is $\gamma$-close to $K_{r}(an)$, $G[U_{i}\setminus Z^{\prime\prime},U_{j}\setminus Z^{\prime\prime}]$ is almost complete for all $i\neq j$. This completes the proof of Theorem 4 with $Z:=Z^{\prime\prime}$. ∎
By (4.9) and (4.11), we get
$f_{1}\geq z_{t}(n)-\varepsilon n^{2-1/t}.$ (4.12)
@@
Now we can give a much cleaner structure, shown in a series of claims below. A key step is to show that $Z^{\prime\prime}\cup W^{\prime\prime}=\emptyset$. From now on we only consider $k\leq 2r$.
Claim 4.7.
Suppose $v_{0}\in V(G)$ and $i\in[r]$ satisfy that $v_{0}$ has at least $\varepsilon n$ neighbors in $Z_{j}$ for every $j\neq i$. Then $v_{0}$ has less than $\varepsilon n$ neighbors in $U_{i}$. In particular, we have $Z^{\prime\prime}=\emptyset$ and $W^{\prime\prime}=\emptyset$.
Proof.
The second part of the claim follows immediately from the definitions of $Z^{\prime\prime}$ and $W^{\prime\prime}$.
Suppose to the contrary, that there exist $v_{0}\in V(G)$ and $i\in[r]$ such that $v_{0}$ has at least $\varepsilon n$ neighbors in $Z_{j}$ for every $j\neq i$ and at least $\varepsilon n$ neighbors in $U_{i}$. Since $|Z_{j}^{j}|\geq(1-\sqrt{\gamma})an$ for all $j\in[r]$, there exist sets $N_{1},\dots,N_{r-1}$ each of size $\varepsilon n-\sqrt{\gamma}n$ such that $N_{j}\subseteq Z_{j}^{j}\cap N(v_{0})$ for $j\neq i$ and $N_{i}\subseteq(Z_{i}^{i}\cup W_{i})\cap N(v_{0})$. Recall that $W_{1}^{\prime}=W_{1}^{\prime}\cap V_{q_{1}}$. By averaging, there exists $N_{1}^{\prime}\subseteq N_{1}$ with $|N_{1}^{\prime}|\geq(\varepsilon n-\sqrt{\gamma}n-2r\gamma n)/2\geq\varepsilon n/3$ such that all vertices of $N_{1}^{\prime}$ are in $Z_{1}^{1}\cup W_{1}^{\prime}$ and from the same cluster, that is,
$N_{1}^{\prime}\subseteq Q,\text{ where }Q\in\{Z_{1}^{1},W_{1}^{\prime}\}.$
Note that $N_{1}^{\prime}\subseteq W_{1}^{\prime}$ is possible only if $i=1$ and $a=1$. If $i\neq 1$, then let $N_{i}^{\prime}:=N_{i}\setminus((W_{i}\setminus W_{i}^{\prime})\cup V_{q_{1}})$ and for every $j\in[r]\setminus\{1,i\}$, let $N_{j}^{\prime}:=N_{j}$. By $(P1)$, $|W_{i}\setminus W_{i}^{\prime}|\leq 2r\gamma n$, and by (4.16), $|W_{i}\cap V_{q_{1}}|\leq\gamma n$. Thus, we have $|N_{j}^{\prime}|\geq\varepsilon n/3$ for all $j\in[r]$. Because the sets $N_{j}^{\prime}$ are small, we can not apply the degree conditions (Deg) to them and instead, we use (D0).
Recall that $B_{1}$ is given by (D0). Next we show that $G[\tilde{U}_{1}\setminus B_{1}]$ does not contain a copy of $K_{t-1,t}$ such that the part of size $t$ is in $N_{1}^{\prime}$. Suppose instead, there is such a copy of $K_{t-1,t}$, with parts denoted by $A$ and $B$, such that $|A|=t$, $A\subseteq N_{1}^{\prime}\setminus B_{1}$ and $B\subseteq\tilde{U}_{1}\setminus B_{1}$. Recall that $N_{i}^{\prime}\cap V_{q_{1}}=\emptyset$ and for each $j\in[r]\setminus\{1,i\}$, $N_{j}^{\prime}\subseteq Z_{j}^{j}$. Observe that for every $v\in\tilde{U}_{1}\setminus B_{1}$, we have $d(v,N_{j}^{\prime})\geq|N_{j}^{\prime}|-\sqrt{\gamma}n$. Indeed, if $j\neq i$, then $N_{j}^{\prime}\subseteq Z_{j}^{j}$ and we have $d(v,N_{j}^{\prime})\geq|N_{j}^{\prime}|-\sqrt{\gamma}n$ by (D0); otherwise note that $N_{i}^{\prime}\subseteq Z_{i}^{i}\cup(W_{i}^{\prime}\cap W_{i})$, and by (D0) and $N_{i}^{\prime}\cap V_{q_{1}}=\emptyset$ we have $d(v,N_{i}^{\prime})\geq|N_{i}^{\prime}|-\sqrt{\gamma}n$. Therefore, we obtain that the vertices in $A\cup B$ have at least $|N_{j}^{\prime}|-(2t-1)\sqrt{\gamma}n\geq(1-\gamma^{1/3})|N_{j}^{\prime}|$ common neighbors in each $N_{j}^{\prime}$, $j\in[2,r]$. Because each pair $N_{j}^{\prime},N_{j^{\prime}}^{\prime}$ has a high density, we can find a copy of $K_{r-1}{(t)}$ in the union of these common neighborhoods, which together with $A\cup B\cup\{v_{0}\}$ form a copy of $K_{r+1}(t)$, a contradiction.
Now given that $G[\tilde{U}_{1}\setminus B_{1}]$ does not contain a copy of $K_{t-1,t}$ such that the part of size $t$ is in $N_{1}^{\prime}\setminus B_{1}$, we give a refined estimate on $f_{1}$. Indeed, since $G[N_{1}^{\prime}\setminus B_{1},Z_{1}^{1}\setminus B_{1}]$ does not contain a copy of $K_{t-1,t}$ such that the part of size $t$ is in $N_{1}^{\prime}\setminus B_{1}$, we get $e_{G}(N_{1}^{\prime}\setminus B_{1},Z_{1}^{1}\setminus B_{1})=O(n^{2-1/(t-1)})$. Similarly $e_{G}(N_{1}^{\prime}\setminus B_{1},W_{1}^{\prime}\setminus B_{1})=O(n^{2-1/(t-1)})$. Suppose $N_{1}^{\prime}\subseteq V_{q}$ for some $q\in\{1,q_{1}\}$, then we have
Recall that $|N_{1}^{\prime}|\geq\varepsilon n/3$ and $|B_{1}|\leq 2\sqrt{\gamma}n$. Therefore, (regardless of $a=1$ or $(a,b)=(2,0)$) we can bound $f_{1}\leq|E(G[\tilde{U}_{1}])|$ by
$f_{1}\leq z_{t}\left((1-\tfrac{\varepsilon}{3})n,n\right)+O(n^{2-1/(t-1)})+O(\sqrt{\gamma}n^{2-1/t})<z_{t}(n)-3rC_{0}kn,$
When $a=2$ and $b=0$ (i.e., $k=2r$), since $Z^{\prime\prime}_{i}=\emptyset$ and $W_{i}=\emptyset$ for all $i\in[r]$, by (4.3), we get $\tilde{U}_{i}=Z_{i}$ for all $i\in[r]$. Therefore $e(G)=e(G^{\prime})+\sum_{i=1}^{r}e_{G}(\tilde{U}_{i})+e(Z^{\prime\prime}\cup W^{\prime\prime},G)\leq t_{r}(k)n^{2}+z_{t}(n)+(r-1)(t-1)n$ by $(K1)$ and (4.11), proving Theorem 3 for $k=2r$.
For the remaining of the proof, we only need to consider $a=1$ (and thus $b>0$). Moreover, now for $i,j\in[r]$ each $Z_{i}^{j}\subseteq Z_{i}$ is an independent set and thus $e_{G}(Z_{i}^{j})=0$. So we can first update our bounds on $e(G)$ and $f_{1}$. Recall the bounds (4.7), (4.8) and (4.11) and we have
$e(G)$ $\leq t_{r}(k)n^{2}+\sum_{i\in[r]}(f_{i}+\beta_{i})$
$\leq t_{r}(k)n^{2}+f_{1}+(r-1)(t-1)(1+\sqrt{\gamma})n+2(r-1)rC_{0}\sqrt{\gamma}n,$
yielding
$f_{1}\geq z_{t}(n,n)-C_{0}n$ (4.14)